OpenStudy (anonymous):

jack usually mows his lawn in 6 hrs. mary mows the same yard in 3 hrs. how much time would it take for them to mow the lawn together.

OpenStudy (saifoo.khan):

2

OpenStudy (anonymous):

see thats what we thought but in thee book is tries to set up some big equation

OpenStudy (saifoo.khan):

the equation will be, \[\frac16+\frac13 = \frac 1x\]

OpenStudy (anonymous):

but it would be 1/2 not 2 right?

OpenStudy (saifoo.khan):

1/2 = 1/x invert both will give u 2.